OpenStudy (anonymous):

How many dimes are there in 4x-1 cents?

OpenStudy (anonymous):

there should be more info to this problem

OpenStudy (anonymous):

Answer choices : 40x-10 , 2/5x-1/10, 40x-1, 4x-1 and 20x-5

OpenStudy (anonymous):

I guess A. 40X-10 but that is a strange question

OpenStudy (anonymous):

How did you get A?

OpenStudy (anonymous):

well a dime is worth 10 cents so youll take the equation 4x-1 and multiply it all by 10 giving you 40x-10

OpenStudy (zarkon):

how many dimes are in 20 cents ...I know I certainly wouldn't multiply by 10 since the answer is 2

OpenStudy (anonymous):

So don't you have to devide that by 100?

OpenStudy (zarkon):

divide by 10

OpenStudy (anonymous):

4x-1 ?

OpenStudy (zarkon):

50 cents ....50/10=5 dimes 70 cents ... 70/10=7 dimes 4x-1 cents ... (4x-1)/10=4/10x-1/10=2/5x-1/10 dimes

OpenStudy (anonymous):

Ah that makes more sense! thank you :D